# Beyond Attraction: A Novel Approach to Repulsive Casimir-Lifshitz Forces using Controllable Off-stoichiometry in Gapped Metals

### v2

@@ IV Conclusions
Here, we have reported the development of a theoretical model that predicts how moving across regions with different off-stoichiometry can be used to tailor the sign of the Casimir-Lifshitz force between the Teflon nanosurface and gapped metal surfaces separated by methanol. Specifically, considering the range of gapped metals Ca6-xAl7O16, we demonstrate that controllable off-stoichiometry regions can be used to tune the Lifshitz interaction from attraction to repulsion. This behavior originates from the effect of off-stoichiometry on the dielectric function of gapped metals - the formation of cation vacancies induces a reduction of free carriers and the Drude contribution to the dielectric function. We show that the phase transition from insulator (Ca5.5Al7O16) to metal (Ca6Al7O16 and Ca5.75Al7O16) leads to a change in the force from attraction to repulsion at the nano to sub-micron scale driven by retardation and caused by the difference in dielectric functions of the gapped metals. The key point in our work is that a zero-frequency contribution and its interconnection with retardation and closeness of dielectric function of the liquids media and PTFE are the main origin for the sign change switching at unprecedented small distances. To the best of our knowledge, in the past purely zero-frequency effects have always been assumed to occur at much larger separations or at much higher temperatures Ninham and Daicic (1998). Here, we see its domination for separations as small as a few nanometers in the metallic case but not for the insulating phase. While the dielectric functions used here do not have any temperature dependence (and hence can’t be used to represent the real temperature-dependent picture), one can find that, e.g., 10% changes in temperature, and the corresponding changes in the zero-frequency term, could lead to a measurable change in attraction-repulsion transition distances. With a change in background liquid media, it is also possible to have a zero-frequency Casimir-Lifshitz term leading to quantum levitation. In this way, using different liquids, we propose a new tool for nanoengineering via phase-transition-related quantum levitation in a liquid medium.
Acknowledgements.
@@
SP, OIM, and MB’s research contributions are part of the project No. 2022/47/P/ST3/01236 co-funded by the National Science Centre and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 945339. The research by SP and MB took place at the “ENSEMBLE3 - Centre of Excellence for nanophotonics, advanced materials and novel crystal growth-based technologies" project (grant agreement No. MAB/2020/14) carried out within the International Research Agendas programme of the Foundation for Polish Science co-financed by the European Union under the European Regional Development Fund, the European Union’s Horizon 2020 research and innovation programme Teaming for Excellence (grant agreement. No. 857543) for support of this work. S.O. thanks the National Science Centre, Poland (grant no. UMO/2020/39/I/ST4/01446) and the “Excellence Initiative - Research University" (IDUB) Program, Action I.3.3 - “Establishment of the Institute for Advanced Studies (IAS)" for funding (grant no. UW/IDUB/2020/25). We gratefully acknowledge Poland’s high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016228 and for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through grant no. PLL/2023/4/016319. M.B. gratefully acknowledges the kind hospitality at the University of Cagliari where this work was completed.

### v3

@@ 4 Conclusions
Here, we have predicted that how moving across regions with different off-stoichiometry can be used to tailor the sign of the Casimir-Lifshitz force between the Teflon nanosurface and gapped metal surfaces separated by methanol. Specifically, considering the range of gapped metals Ca6-xAl7O16, we demonstrate that controllable off-stoichiometry regions can be used to tune the Lifshitz interaction from attraction to repulsion. This behavior originates from the effect of off-stoichiometry on the dielectric function of gapped metals - the formation of cation vacancies induces a reduction of free carriers and the Drude contribution to the dielectric function. We show that the phase transition from insulator (Ca5.5Al7O16) to metal (Ca6Al7O16 and Ca5.75Al7O16) leads to a change in the force from attraction to repulsion at the nano to sub-micron scale driven by retardation and caused by the difference in dielectric functions of the gapped metals. The key point in our work is that a zero-frequency contribution and its interconnection with retardation and closeness of dielectric function of the liquids media and PTFE are the main origin for the sign change switching at unprecedented small distances. To the best of our knowledge, in the past purely zero-frequency effects have always been assumed to occur at much larger separations or at much higher temperatures [43]. Here, we see its domination for separations as small as a few nanometers in the metallic case but not for the insulating phase. While the dielectric functions used here do not have any temperature dependence (and hence can’t be used to represent the real temperature-dependent picture), one can find that, e.g., 10% changes in temperature, and the corresponding changes in the zero-frequency term, could lead to a measurable change in attraction-repulsion transition distances. With a change in background liquid media, it is also possible to have a zero-frequency Casimir-Lifshitz term leading to quantum levitation. In this way, using different liquids, we propose a new tool for nanoengineering via phase-transition-related quantum levitation in a liquid medium.
CRediT authorship contribution statement
Subhojit Pal: Writing – review -& editing, Writing – original draft, Software development, Investigation. Silvio Osella: Writing – review & editing. Oleksandr Malyi: Writing – review & editing, Materials modeling, Supervised computations, Computational resources acquisition. Mathias Boström: Writing – review -& editing, Writing – original draft, Initiated and supervised the project, Funding acquisition.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgement
@@
SP, OIM, and MB’s research contributions are part of the project No. 2022/47/P/ST3/01236 co-funded by the National Science Centre and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 945339. Institutional and infrastructural support for the ENSEMBLE3 Centre of Excellence was provided through the ENSEMBLE3 project (MAB/2020/14) delivered within the Foundation for Polish Science International Research Agenda Programme and cofinanced by the European Regional Development Fund and the Horizon 2020 Teaming for Excellence initiative (Grant Agreement No. 857543), as well as the Ministry of Education and Science initiative “Support for Centres of Excellence in Poland under Horizon 2020” (MEiN/2023/DIR/3797). S.O. thanks the National Science Centre, Poland (grant no. UMO/2020/39/I/ST4/01446) and the “Excellence Initiative - Research University” (IDUB) Program, Action I.3.3 - “Establishment of the Institute for Advanced Studies (IAS)” for funding (grant no. UW/IDUB/2020/25). We gratefully acknowledge Poland’s high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016228 and for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through grant no. PLL/2023/4/016319. M.B. gratefully acknowledges the kind hospitality at the University of Cagliari where this work was completed.